# KnowGPT: Knowledge Injection for Large Language Models

### v4

@@ 5.2.1 Leaderboard Ranking
$\mathcal{P}_{\text{sub}}$ ChatGPT 0.750 0.739 0.865 0.695
$\mathcal{P}_{\text{RL}}$ ChatGPT 0.815 0.800 0.889 0.755
@@
Table 4: Ablation study on different prompt translation formats for the extracted knowledge.
Path Extraction Prompts CSQA OBQA MedQA
@@
5.3 Ablation Studies (RQ2)
To answer RQ2, we conduct two ablation studies. First, in Table 3, we measure the importance of the tailored reinforcement learning-based path extraction module, i.e., $\mathcal{P}_{\text{RL}}$. Specifically, we compare it with the heuristic sub-graph extraction strategy, i.e., $\mathcal{P}_{\text{sub}}$. The performance is evaluated by directly feeding the extracted knowledge with the prompt format of ‘Sentence’, i.e., $\mathcal{F}_{s}$, to ChatGPT. We also include ‘w/o KG’ as the baseline where ChatGPT is asked to independently answer the given question with no reasoning background provided. The results clearly indicate the vital role of our proposed path extraction strategies. Second, we compare each of the three prompt formats subject to the same extracted knowledge. The detailed results are shown in Table 4. Though different formats perform similarly within the difference of 2.2% - 3.3%, they are particularly suitable for different kinds of questions. We illustrate this observation in the following case study section. Both ablation studies support the indispensability of each module, armed with a tailored deep reinforcement learning-based path extraction and a context-aware prompt translation, our KnowGPT performs best on all three benchmark datasets.
6 Conclusion and Future Work
In this work, we formally define the problem of knowledge injection for LLMs in complex question answering. A novel framework, namely KnowGPT, is presented to integrate KGs into LLMs effectively with model APIs only. We first train a deep RL policy to extract informative and concise reasoning paths from the KG. Then we learn an MAB to select the most suitable and effective path extraction method and prompt template subject to each question. Extensive experiments on both general and domain-specific QA datasets show the superior performance of KnowGPT and the effectiveness of each component. In the future, we will study more advanced path extraction strategies and prompt templates to improve the performance further.
Limitations
@@ 6 Conclusion and Future Work
Through our exploration, we realize the natural limitations of KnowGPT brought by real-world KGs. Existing KGs are automatically constructed based on online corpora. This inevitably introduces a considerable number of noisy triples into KGs. Under such circumstances, the noisy knowledge may mislead the LLMs to wrong predictions despite the effectiveness of our injection methods. To apply KnowGPT to practical scenarios, we would leverage off-the-shelf KG refinement algorithms to improve the quality of KGs.
Ethics Statement
In this study, all the datasets are publicly available and have been extensively used in research related to neural language processing.
Appendix A Case Studies (RQ3)
For RQ3, we provide insights into how KnowGPT facilitates the prompt translation with a real case from CommonsenseQA. We visualize both the extracted knowledge and the textual inputs to ChatGPT in Figure 3. In this example, given the same extracted knowledge, ChatGPT answers correctly based on the sentence format that we provide. In contrast, it fails to answer the question with triples and graph descriptions. They clearly indicate the superiority of KnowGPT in an automatic context-aware prompt translation. We make the following observations: $(i)$ Triple format $\mathcal{F}_{t}$ is intuitively suitable for all the simple questions by directly indicating the one-hop knowledge. $(ii)$ Graph description may inevitably introduce noise to ensure the completeness and contextual fluency of the directed graph. In this example, since ‘vacation’ appears in both question and answer choices, over-emphasizing and connecting the knowledge about ‘vacation’ with other concepts in the graph misleads the model to make a prediction with an oblique focus. $(iii)$ Our KnowGPT has shown superior performance in automatically constructing suitable prompts for particular questions.
@@ Appendix B Implementation Details
B.1 Entity Linking and Heuristic Path Extraction $\mathcal{P}_{sub}$
For each QA context, we adopt the methodology outlined in the prior research Lin et al. (2019); Yasunaga et al. (2022) to extract the subgraph from the background knowledge graph (KG), denoted as $\mathcal{G}$. We commence by executing entity linking on $\mathcal{G}$, resulting in an initial collection of nodes, $V_{topic}$. Next, we incorporate bridge entities that appear within a 2-hop path between any two linked entities from $V_{topic}$, yielding the set $V_{retrieval}$. Subsequently, we refine this set by evaluating the relevance score for each node, adhering to the method proposed Yasunaga et al. (2022). From this refined set, only the top 200 nodes, based on score, are retained, and the others are discarded. We then extract all edges connecting any pair of nodes in $V_{sub}$, creating the retrieved subgraph $G_{sub}$. Each node within $G_{sub}$ is designated a type based on its association to either the topic entities $Q$ or target entities $A$. Intuitively, the relevant reasoning background lies in a question-specific subgraph $\mathcal{G}_{sub}$ that contains all the source entities $S$, target entities $A$, and their $k$-hop neighbors. Therefore, the reasoning background could be provided as the $\mathcal{G}_{sub}$, we denote this direct path extraction method as $\mathcal{P}_{sub}$.
B.2 Graph Initialization
To calculate the initial node embeddings for entities extracted from the background KG, we adopt the approach proposed by the previous study Feng et al. (2020). Specifically, we transform knowledge triples from the KG into sentences and feed them into pre-trained LMs to get node embeddings. Specifically, to ensure a fair comparison, we implement all the KG-enhanced baselines and our model with the same advanced language models that are optimized for particular datasets. Specifically, RoBert-large Liu et al. (2019) is used for CommenseQA, while AristoRoBERTa Clark et al. (2020) is designated for OpenBookQA. For MedQA, we opt for the top-tier biomedical language model, SapBERT Liu et al. (2021), to enhance comprehension of the biomedical field.
B.3 Datasets
We evaluate KnowGPT on three QA datasets spanning two fields: CommonsenseQA Talmor et al. (2019) and OpenBookQA Mihaylov et al. (2018) serve as benchmarks for commonsense reasoning, while MedQA-USMLE Jin et al. (2021) acts as a domain-specific QA benchmark. While the official test set serves primarily for leaderboard rankings, we initially assess model efficacy using the in-house (IH) data split introduced in Lin et al. (2019). The official dataset is denoted as CSQA, while the IH split is represented by CSQA(IH)*. The statistics of these three datasets can be found in Table 5.
@@ B.3 Datasets
CommonsenseQA is a multiple-choice question-answering dataset, each question accompanied by five potential answers. Answering its 12,102 questions necessitates a foundation in commonsense knowledge. While the official test set serves primarily for leaderboard rankings, we initially assess model efficacy using the in-house (IH) data split introduced in Lin et al. (2019). The official dataset is denoted as CSQA, while the IH split is represented by CSQA(IH)*.
@@
MedQA-USMLE, abbreviated as MedQA, is a dataset consisting of 4-option multiple-choice questions that demand a grasp of biomedical and clinical understanding. These questions are sourced from preparatory tests for the United States Medical Licensing Examinations, and the dataset encompasses 12,723 questions. We adhere to the original data divisions as outlined in Jin et al. (2021).
Background Knowledge To facilitate common sense reasoning, we employ ConceptNet Speer et al. (2017), an extensive commonsense knowledge graph comprising more than 8 million interconnected entities through 34 concise relationships. For tasks specific to the medical domain, we leverage USMLE Yasunaga et al. (2021) as our foundational knowledge source. USMLE is a biomedical knowledge graph that amalgamates the Disease Database segment of the Unified Medical Language System (UMLS) Bodenreider (2004) and DrugBank Wishart et al. (2017). This repository encompasses 9,958 nodes and 44,561 edges.
B.4 Implementation of Baselines
To verify the effectiveness of our proposed KnowGPT, we carefully selected baseline models from three aspects to ensure a comprehensive evaluation, among which Bert-base, Bert-large Kenton and Toutanova (2019), and RoBert-large Liu et al. (2019) are picked for being representative fine-tune LM methods; MHGRN Feng et al. (2020), QA-GNN Yasunaga et al. (2021), HamQA Dong et al. (2023a), JointLK Sun et al. (2022), GreaseLM Zhang et al. (2022b) and GrapeQA Taunk et al. (2023) represent the state-of-art KG-enhanced LMs; ChatGLM Du et al. (2022), ChatGLM2 Zeng et al. (2022), Baichuan-7B, InternLM Team (2023), GPT-3 Brown et al. (2020), ChatGPT Ouyang et al. (2022) and GPT-4 OpenAI (2023) are picked for being representative generative large language models. Notably, while some LLM baselines are actually open-source, we conduct the question-answering task under the zero-shot setting with the question query from the test set as input. All baseline methods used in this paper are based on their open-source implementations or officially-released APIs. We list the source links in Table 6. Notably, we used the model ‘text-davinci-002’ provided by OpenAI as the implementation of GPT-3, and ‘gpt-3.5-turbo’ and ‘gpt-4’ as the implementations of ChatGPT and GPT-4, respectively.
Table 6: Implementation codes for baselines.
Appendix C Subplementary Results
C.1 Fine-tune LLMs
Table 7: Fine-tuned LLMs on CommonsenseQA.
LLM Acc
ChatGLM 0.559
ChatGLM2 0.600
LLaMA-7B 0.650
Baichuan-7B 0.588
Alpaca-7B 0.687
Vicuna-7B 0.667
InternLM-7B 0.752
KnowGPT 0.818
C.2 The effect of prompt format on different types of questions.
@@ C.1 Fine-tune LLMs
$\mathcal{F}_{t}$ (Triple) 0.9412 0.7447 0.4210
$\mathcal{F}_{s}$ (Sentence) 0.8823 0.8298 0.4739
$\mathcal{F}_{g}$ (Graph) 0.8529 0.7021 0.7895

### v5

@@ 4.3 Ablation Studies (RQ2)
$\mathcal{P}_{\text{sub}}$ GPT-3.5 0.750 0.739 0.865 0.695
$\mathcal{P}_{\text{RL}}$ GPT-3.5 0.815 0.800 0.889 0.755
@@
Table 4: Ablation study on different prompt formats for the extracted knowledge.
Knowledge Extraction Prompts CSQA OBQA MedQA
@@
5 Limitation
@@
Through our exploration, we realize the natural limitations of KnowGPT brought by real-world KGs. Existing KGs are automatically constructed based on online corpora. This inevitably introduces a considerable number of noisy triples into KGs. The noisy knowledge may mislead the LLMs to wrong predictions despite the effectiveness of our prompting methods. In the future, we would leverage KnowGPT with off-the-shelf KG refinement algorithms to improve the quality of KGs.
6 Conclusion
The main objective of this paper is to tackle the hallucination issue that arises when applying LLM to specific domains. Although LLMs have strong reasoning capabilities, they still struggle to answer professional questions in specific domains, especially when the pre-training corpus lacks relevant knowledge. To address this issue, we propose a KG-augmented LLM model, named KnowGPT, which injects relevant domain knowledge from KGs into LLMs to assist the LLM in accurately answering professional questions. A novel framework, namely KnowGPT, is presented to integrate KGs into LLMs effectively with model APIs only. We first train a deep RL policy to extract informative and concise reasoning background from the KG. Then we learn an MAB to select the most effective knowledge extraction method and prompt template for each question. Extensive experiments on both general and domain-specific QA show superior performance of KnowGPT compared to all competitors.
Appendix A Related Work
To address the hallucination issues that arise when applying LLMs to specific domains, researchers have proposed various methods, which can be broadly categorized into three main categories: Fine-Tuning (FT), Retrieval-Augmented Generation (RAG) and KG-enhanced LLMs.
A.1 Fine-tune LLMs with Domain-specific Data
To help the language model generate more accurate responses in specific domains, extensive studies have been conducted to investigate the efficacy of fine-tuning LLMs with specialized data [19, 28, 29, 15]. Fine-tuning involves training the pre-trained LLM on a smaller dataset of domain-specific text, allowing the model to adapt its knowledge to the target domain. By exposing the LLM to domain-specific vocabulary, terminology, and patterns, fine-tuning enables the model to generate more accurate and relevant responses for domain-specific tasks [19]. Fine-tuning LLMs with domain-specific data has been successfully applied across various domains. For example, in the healthcare domain, fine-tuned LLMs have been used for clinical note analysis [1], biomedical text mining [29], and medical dialogue [58]. Similarly, in the legal domain, fine-tuned LLMs have shown promise in tasks such as legal document classification [6], contract analysis [7], and legal judgment prediction [77].
However, recent research has highlighted the limitations and risks associated with fine-tuning LLMs using domain-specific data [26]. A study conducted by Google Research highlighted that using fine-tuning to update LLMs’ knowledge can be problematic, particularly when the new knowledge conflicts with pre-existing information [16]. In such cases, acquiring new data through FT can lead to the model generating new hallucinations and even experiencing catastrophic forgetting [26, 67]. Besides that, many state-of-the-art LLMs are confined to a black-box role in practice. For instance, Calaude3 [3] and GPT-4 [43] exclusively grant access through their APIs, which means we can only retrieve model responses by submitting textual inputs, with model specifics inaccessible. This lack of access to model architecture and parameters prevents us from employing these fine-tuning methods.
A.2 Retrieval-Augmented Generation for LLMs
Recently, Retrieval-Augmented Generation (RAG) models have been extensively explored to enhance LLMs with external knowledge from text corpora or online sources [30, 76, 42]. Combining LLMs with external knowledge retrieval systems can help reduce hallucinations. However, these approaches face challenges in domain-specific applications: $(i)$ Data quality. Domain knowledge is often scattered across different sources, such as textbooks, research papers, technical manuals, and industry reports [31]. These textual documents may have varying levels of quality, accuracy, and completeness, leading to potential inconsistencies or errors in the retrieved knowledge [78]. $(ii)$ Knowledge hierarchy. Domain knowledge is always complex and hierarchical and contains a high degree of jargon, acronyms, and specialized terminology. However, textual documents typically present this information in a linear and unstructured manner. The lack of explicit relationships and structured organization limits the reasoning and inference capabilities of RAG models, as they cannot leverage the structured connections within domain knowledge to derive new insights or generate more contextually appropriate responses [47]. $(iii)$ Huge search space. There is a lot of domain-irrelevant information in these textual documents, while domain-specific terminologies are always sparsely distributed over these documents [50]. The retrieval model can be computationally expensive and time-consuming, especially when dealing with large-scale knowledge sources, as the model needs to search through vast amounts of unstructured text to find relevant information [40].
A.3 Integration of KGs and LLMs
KG-enhanced LLM aims to leverage the structured knowledge in knowledge graphs [32, 33, 49, 72, 73] to ground the model’s responses in established facts [20, 66, 45]. This grounding ensures that the generated outputs are based on reliable information rather than arbitrary or fabricated statements.
Integrating KGs during Training. Earlier studies adopted a heuristic way to inject knowledge from KGs into the LLMs during pre-training or fine-tuning. For example, ERNIE [52] incorporates entity embeddings and aligns them with word embeddings in the pre-training phase, encouraging the model to better understand and reason over entities. KnowBERT [46] integrates entity linkers with BERT to introduce knowledge about entities during fine-tuning for downstream tasks.
Integrating KGs during Inference. Another line of work focuses on retrieving relevant knowledge from KGs at inference time to augment the language model’s context. Typically, K-BERT [37] uses an attention mechanism to select relevant triples from a KG based on the input context, which are then appended to the input sequence. Similarly, KEPLER [62] learns joint embeddings of text and KG entities, enabling efficient retrieval of relevant knowledge to enhance the model’s predictions. Later works further integrated graph neural networks alongside LLMs for joint reasoning [69, 74, 10] and introduced interactions between text tokens and KG entities within the intermediate layers of LLMs [53, 56].
KG Prompting for LLM. The methods above all assume they know everything about LLMs, including model architectures and parameters. However, as LLMs have been keeping evolving, many SOTA LLMs are confined to a black-box role in practice. As such, the research focus has shifted towards KG prompting that enhances fixed LLMs with KG-based hard prompts. KG Prompting for LLMs has been a new learning paradigm in natural language processing [36, 45]. Specifically, CoK [60] introduces a novel Chain-of-Knowledge prompting to decompose LLM-generated reasoning chains into evidence triples. It then verifies the triples’ factuality and faithfulness using an external knowledge graph, ensuring the accuracy and reliability of the LLM outputs. By allowing the models to comprehend and reason over structured KG inputs, Mindmap [63] enhances the LLMs’ ability to incorporate external knowledge, reduce hallucinations, and provide more transparency into their decision-making process. RoG [39] presents a planning-retrieval-reasoning framework that synergizes the strengths of LLMs and KGs to enhance the reasoning capabilities of language models in a more transparent and interpretable manner. Instead of retrieving factual information from KGs, KGR [18] proposes to autonomously retrofit the initial draft responses generated by LLMs. This innovative technique leverages the knowledge stored within KGs to mitigate hallucination by effectively extracting, verifying, and refining factual information throughout the entire reasoning process of LLMs.
Our model. Despite the promising performance of existing KG prompting methods, their extensive deployment and practical implementation in domain-specific applications are still impeded by two critical issues. The cost associated with calling the LLM API or deploying LLMs with cloud services is prohibitive. For example, GPT-4 is estimated to cost at least thousands of dollars for pilot-scale customer service, making the careful selection of the most informative triples from KGs essential to minimize costs. LLMs are highly sensitive to prompts, with even minor variations in prompts conveying the same semantic meaning potentially yielding drastically different responses. However, existing methods rely on manually designed or rule-based prompts to present factual knowledge from KGs. These hard prompts are always fixed and rigid, lacking the flexibility to adapt to variations in question semantics and KG structures. In this paper, we shed light on the whole community with a novel prompt learning method that is rather effective, generalizable and cost-efficient. There are two key components in our proposed KnowGPT including (i) knowledge retrieval, which leverages deep RL to extract relevant knowledge from KGs, and (ii) context-aware decision module to translate the structured knowledge from KGs into the appropriate prompt format.
@@ Appendix B Implementation Details
B.1 Entity Linking and Heuristic Knowledge Extraction
For each QA context, we adopt the methodology outlined in the prior research [34, 68] to extract the subgraph from the background knowledge graph (KG), denoted as $\mathcal{G}$. We commence by executing entity linking on $\mathcal{G}$, resulting in an initial collection of nodes, $V_{topic}$. Next, we incorporate bridge entities that appear within a 2-hop path between any two linked entities from $V_{topic}$, yielding the set $V_{retrieval}$. Subsequently, we refine this set by evaluating the relevance score for each node, adhering to the method proposed [68]. From this refined set, only the top 200 nodes, based on score, are retained. We then extract all edges connecting any pair of nodes in $V_{sub}$, creating the retrieved subgraph $G_{sub}$. Each node within $G_{sub}$ is designated a type based on its association to either the topic entities $Q$ or target entities $A$. Intuitively, the relevant reasoning background lies in a question-specific subgraph $\mathcal{G}_{sub}$ that contains all the source entities $S$, target entities $A$, and their $k$-hop neighbors. Therefore, the reasoning background could be provided as the $\mathcal{G}_{sub}$, we denote this direct knowledge extraction method as $\mathcal{P}_{sub}$.
B.2 Feature Initialization of Background KG
To calculate the initial node embeddings for entities extracted from the background KG, we adopt the approach proposed by the previous study [13]. Specifically, we transform knowledge triples from the KG into sentences and feed them into pre-trained LMs to get node embeddings. Specifically, to ensure a fair comparison, we implement all the KG-enhanced baselines and our model with the same advanced language models that are optimized for particular datasets. Specifically, RoBert-large [38] is used for CommenseQA, while AristoRoBERTa [9] is designated for OpenBookQA. For MedQA, we opt for the top-tier biomedical language model, SapBERT [35], to enhance comprehension of the biomedical field.
B.3 Statistical Analysis of Datasets
We evaluate KnowGPT on three QA datasets spanning two fields: CommonsenseQA [55] and OpenBookQA [41] serve as benchmarks for commonsense reasoning, while MedQA-USMLE [23] acts as a domain-specific QA benchmark. While the official test set serves primarily for leaderboard rankings, we initially assess model efficacy using the in-house (IH) data split introduced in [34]. The official dataset is denoted as CSQA, while the IH split is represented by CSQA(IH)*. The statistics of these three datasets can be found in Table 5.
@@ B.3 Statistical Analysis of Datasets
CommonsenseQA is a multiple-choice question-answering dataset, each question accompanied by five potential answers. Answering its 12,102 questions necessitates a foundation in commonsense knowledge. While the official test set serves primarily for leaderboard rankings, we initially assess model efficacy using the in-house (IH) data split introduced in [34]. The official dataset is denoted as CSQA, while the IH split is represented by CSQA(IH)*.
MedQA-USMLE, abbreviated as MedQA, is a dataset consisting of 4-option multiple-choice questions that demand a grasp of biomedical and clinical understanding. These questions are sourced from preparatory tests for the United States Medical Licensing Examinations, and the dataset encompasses 12,723 questions. We adhere to the original data divisions as outlined in [23].
Background Knowledge To facilitate common sense reasoning, we employ ConceptNet [51], an extensive commonsense knowledge graph comprising more than 8 million interconnected entities through 34 concise relationships. For tasks specific to the medical domain, we leverage USMLE [69] as our foundational knowledge source. USMLE is a biomedical knowledge graph that amalgamates the Disease Database segment of the Unified Medical Language System (UMLS) [4] and DrugBank [64]. This repository encompasses 9,958 nodes and 44,561 edges.
B.4 Implementation of Baselines
To verify the effectiveness of our proposed KnowGPT, we carefully selected baseline models from three aspects to ensure a comprehensive evaluation, among which Bert-base, Bert-large [24], and RoBert-large [38] are picked for being representative fine-tune LM methods; MHGRN [13], QA-GNN [69], HamQA [10], JointLK [53], GreaseLM [75] and GrapeQA [56] represent the state-of-art KG-enhanced LMs; ChatGLM [12], ChatGLM2 [70], Baichuan-7B, InternLM [57], GPT-3 [5], GPT-3.5 [44] and GPT-4 [43] are picked for being representative generative large language models. To verify the effectiveness of our proposed prompting strategy used KnowGPT, we also add the state-of-the-art KG prompting methods, i.e., CoK [60], RoG [39], and Mindmap [63] as baselines. Notably, while some LLM baselines are actually open-source, we conduct the question-answering task under the zero-shot setting with the question query from the test set as input. All baseline methods used in this paper are based on their open-source implementations or officially-released APIs. Notably, we used the model ‘text-davinci-002’ provided by OpenAI as the implementation of GPT-3, and ‘gpt-3.5-turbo’ and ‘gpt-4’ as the implementations of GPT-3.5 and GPT-4, respectively. All baseline methods used in this paper are based on their open-source implementations. We list the source links in Table 6. All models are implemented in Pytorch and trained on an RTX 3090 with 24 RAM. We use their best configuration as the default value for other distinctive hyperparameters in the baseline methods. To reduce the randomness, we use the random seed and report the average results of three runs.
Table 6: Implementation codes for baselines.
Appendix C Supplementary Results
C.1 Compared to Fine-tune LLMs
Table 7: Fine-tuned LLMs on three benchmark datasets.
LLM CommonsenseQA OpenBookQA MedQA
ChatGLM 55.9% 54.2% 40.4%
ChatGLM2 60.0% 55.6% 45.5%
LLaMA-7B 65.0% 58.8% 46.9%
Baichuan-7B 58.8% 56.1% 43.4%
Alpaca-7B 68.7% 63.3% 47.1%
Vicuna-7B 66.7% 64.0% 45.2%
InternLM-7B 75.2% 70.7% 48.3%
KnowGPT 81.8% 92.4% 78.1%
C.2 Case Studies (RQ3)
For RQ3, we provide insights into how KnowGPT facilitates the prompt translation with a real case from CommonsenseQA. We visualize both the extracted knowledge and the textual inputs to GPT-3.5 in Figure 3. In this example, given the same extracted knowledge, GPT-3.5 answers correctly based on the sentence format that we provide. In contrast, it fails to answer the question with triples and graph descriptions. They clearly indicate the superiority of KnowGPT in an automatic context-aware prompt translation. We make the following observations: $(i)$ Triple format $\mathcal{F}_{t}$ is intuitively suitable for all the simple questions by directly indicating the one-hop knowledge. $(ii)$ Graph description may inevitably introduce noise to ensure the completeness and contextual fluency of the directed graph. In this example, since ‘vacation’ appears in both question and answer choices, over-emphasizing and connecting the knowledge about ‘vacation’ with other concepts in the graph misleads the model to make a prediction with an oblique focus. $(iii)$ Our KnowGPT has shown superior performance in automatically constructing suitable prompts for particular questions.
@@ C.2 Case Studies (RQ3)
C.3 The effect of prompt format on different types of questions.
Generally, based on the complexity of the reasoning background, questions can be roughly categorized into three classes: (i) Simple question, like “What could be used as an electrical conductor?”. (ii) Multi-hop reasoning question, like “Which school did Bill Gates’ wife graduate from?” (iii) Global reasoning questions: “What do cats have in common with most mammals?”
Different types of questions correspond to different reasoning background. For example, simple questions only require basic factual triples, while multi-hop reasoning questions need a reasoning chain, and global reasoning questions require a more complex tree/graph-like reasoning background. To convert the extracted raw knowledge into textual prompt with corresponding logical reasoning structure, we designed three different prompt templates, including $F_{t}$,$F_{s}$ and $F_{g}$.
To verify the effectiveness of prompt formats, we conducted a statistical analysis on the benchmark datasets in terms of question types and then separately calculated the accuracy of different prompt formats on specific types of questions. As shown in Table 8, we can observe that triple-based prompt performs best on simple questions while graph description-based prompt performs better than any other prompt formats on complex questions. This is because graph description-based prompt could provide LLMs with more detailed and structured information by highlighting the local structure of the central entity.
@@
$\mathcal{F}_{t}$ 94.12% 74.47% 42.10%
$\mathcal{F}_{s}$ 88.23% 82.98% 47.39%
$\mathcal{F}_{g}$ 85.29% 70.21% 78.95%
C.4 The Effect of two Knowledge Retrieval Methods
If the background knowledge graph is complete and high-quality, our proposed RL-based knowledge extraction can ideally handle all cases. But in pratical implementation, we found that there are a few long-tail entities in KGs which only have few neighbors. In such cases, RL-based knowledge extraction can hardly retrieve reachable or effective paths due to the incompletion and sparsity of the background graph. Instead, directly extracting the whole subgraph through rule-based subgraph extraction can provide more knowledge for reasoning. The RL-based retrieval method $\mathcal{P}_{RL}$ could make sure our model extracts more concise and informative knowledge in most cases, while the rule-based subgraph extraction $\mathcal{P}_{RL}$ could supplement the former especially when the reasoning background is sparse. Thus, in this paper, we adopt Multi-Armed Bandit (MAB) to automatically select the most suitable retrieval methods from these two candidates, i.e., $\mathcal{P}_{RL}$ and $\mathcal{P}_{sub}$. The ablation study on the knowledge retrieval method can be found in Table 4 of the main content. Specifically, we compare our model with $\mathcal{P}_{RL}$, the heuristic sub-graph extraction $\mathcal{P}_{sub}$, and we also include ‘w/o KG’ as the baseline where the LLM is asked to independently answer the given question with no reasoning background provided. The results shown in Table 3 of our submission clearly indicate the vital role of our proposed knowledge extraction strategies.
C.5 Efficiency Analysis: API Cost and Model Efficiency
To investigate the efficiency of our model, we compare it on MedQA with six representative baselines from three different categories, including traditional KG-enhanced LMs, i.e., JointLK and GrapQA, zero-shot LLMs, i.e., ChatGPT and GPT-4, and the SOTA KG-prompting-based LLMs, i.e., CoK and Mindmap. Notably, zero-shot LLMs, like ChatGPT and GPT-4, do not require training while traditional KG-enhanced LM methods like JointLK and GrapeQA are based on lightweight LMs and do not require interaction with LLMs, thus they do not incur API cost.
From Table 9, we can see that (i) Traditional KG-enhanced LM methods, like JointLK and GrapQA, have the shortest inference time. This is because the other models need to send requests to the Black-box LLM via API, and the extra response time of the LLM leads to long inference times for these models. Despite the efficiency, they have the worst performance. (ii) Our model outperforms the other models with comparable training time and the most economical API cost compared to models in the same category, including CoK and Mindmap. That is because we are the only model that consider the conciseness of knowledge during retrieval and prompt design.
Table 9: API cost and model efficiency analysis on MedQA.
Model Training time Inference time Avg tokens Cost(\$) Performance
JointLK 4.02 h 0.13 h - - 40.3%
GrapeQA 4.49 h 0.15 h - - 39.5%
ChatGPT - 0.24 h 98 1.87 48.7%
GPT-4 - 0.21 h 98 29.77 76.3%
CoK 5.85 h 0.44 h 1129 21.54 72.2%
Mindmap 4.93 h 0.36 h 761 14.52 75.1%
KnowGPT 5.47 h 0.33 h 348 6.64 78.1%
C.6 Stability Analysis
Due to page limits, we carefully described how we designed the reward function to ensure the RL model captures effective and informative knowledge in the main content of our submission. Actually, we have also employed the following techniques to stabilize the training process.
Gradient clipping for policy network. The policy network is updated using gradient-based optimization to maximize the expected cumulative reward. To stabilize the training of the policy network, we applied gradient clipping [14] to prevent the gradients from becoming too large. This technique helps mitigate the problem of exploding gradients and ensures stable updates of the network parameters. To demonstrate the effectiveness of gradient clipping, we conducted additional experiments comparing the performance of our model with and without gradient clipping across five runs. The results are presented in the table below. As shown in Table 11, the model with gradient clipping consistently achieves better performance across all datasets, demonstrating the stability and robustness of our method.
Exploration-exploitation balance for MAB. In our designed MAB, we have introduced a penalty term, i.e., $\beta$, to encourage exploration on under-explored arms (i.e., prompt formats), otherwise, it will lead to a greedy selection based on $c\times\alpha_{(i)}$. Table 11 shows the performance comparison between with and without this term.
Table 10: Stability study on gradient clipping.
Model Variant CSQA OBQA MedQA
w. gradient clipping 81.8 $\pm$ 0.51 92.4 $\pm$ 0.69 78.1 $\pm$ 0.66
w/o gradient clipping 78.9 $\pm$ 1.92 90.2 $\pm$ 1.25 75.5 $\pm$ 2.01
Table 11: Stability analysis on penalty term.
Model Variant CSQA OBQA MedQA
w. penalty term 81.8 $\pm$ 0.51 92.4 $\pm$ 0.69 78.1 $\pm$ 0.66
w/o penalty term 79.5 $\pm$ 1.63 91.0 $\pm$ 1.45 76.4 $\pm$ 1.89
Appendix D Broader Impacts
We are dedicated to not only focusing on the specific NLP task, i.e., knowledge-based QA, but also inspiring and benefitting a wide range of NLP communities with a novel knowledge injection framework to fast adapt LLMs to specific domains. Specifically, we would like to declare the following two contributions for a wide range of NLP communities.
Exploring Effective Knowledge Injection for LLMs: Applying LLMs for various NLP-related tasks in different domains is gaining momentum. However, fine-tuning LLMs for domain knowledge-based QA is expensive. Our paper presents a remarkably strong framework that could (i) retrieve the reasoning background from domain knowledge graph and (ii) automatically find out the optimal prompt to inject the domain knowledge into the LLM.
Providing Insights in Prompt Learning: While hard prompt is still dominating prompt learning for the majority of LLM-based scenarios, we shed light on the community with a novel prompt learning method that is rather generalizable and effective. We propose an auto-selection algorithm that can determine the best prompt format for LLMs to maximally understand the reasoning background.